Gifts and More (G&M) is a small private business that deals in gift items. The company has $18,000 in working capital requirements and fixed assets with a fair value of $190,000. G&M’s required returns on working capital, fixed assets, and intangible assets are estimated as 2.5%, 8%, and 12%, respectively. Normalized earnings for the year just ended amounted to $91,000 and residual income is expected to grow at 3% per year. The company's excess earnings are closest to: Select one:

a. $84,810

b. $75,350

c. $81,650
Answer #1

Working Capital = 18000

Return on working capital = 18000*2.5% = 450

Fair value of fixed assets = 190000

Return on fixed assets = 190000*8% = 15200

Net return on total assets = 450+15200 = 15650

Normal earnings = 91000

company's excess earnings = Normal earnings - Net return on total assets

= 91000 - 15650 = 75350

Answer is b. $75,350